# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ 15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
Table 7: The six task‐sequence orders used in our continual learning experiments. Sequences 1–3 follow the standard CL benchmarks employed in prior work. Sequences 4–6 extend to longer 15‐task streams, as introduced in (Razdaibiedina et al., 2023).
@@
A.1 Detailed Experiment Settings
CL on CLIP. We follow the experimental setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) and use the CLIP model with a ViT-B/16 backbone (Radford et al., 2021) for all experiments. By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update. Each task is trained for 500 iterations using AdamW (Loshchilov & Hutter, 2017) with a learning rate of 5e45e-4. During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task. The rank activation budget is set to 16 throughout all tasks. We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
CL on LMs. We follow the protocol of previous work in continually fine-tuning the T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on a suite of text-classification tasks. We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024). To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark (Zhang et al., 2015), GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and the IMDB movie reviews corpus. Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 7). After completing the final task in each stream, we report the average accuracy across all tasks. All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1. MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in (Wang et al., 2023; Qiao & Mahdavi, 2024). We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) using llama-Factory (Zheng et al., 2024) and evaluate using lm-eval-harness (Gao et al., 2024) on zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model. We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4. We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks. The self-routed gating uses a temperature $\tau_{\text{MoRA}}=0.5$ and a threshold $\delta=0.2$. We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
Further details of the used datasets in MTIL and X-TAIL. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010). In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024). In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
@@ A.1 Detailed Experiment Settings
Table 8: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from (Qiao & Mahdavi, 2024).
@@
MoRA 77.8 78.0 79.3 78.4
A.2 Continual Learning of LMs on Long Task Sequences
In Sec. 4.2 Table 3, we evaluate MoRA on standard CL benchmarks (Razdaibiedina et al., 2023). In Table 8, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in (Wang et al., 2023; Qiao & Mahdavi, 2024). Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning. Two key design choices drive this robust performance in the long‐sequence regime. First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold. At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference. Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows. Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
Table 10: Comparisons of trainable parameters for each pre-trained weight matrix during continual learning of each task. MoE-LoRA and MoE-Adapter trains additional router module with LoRA experts. In MoRA, $k$ denotes the rank activation budget (with $k\leq r$).
@@ A.2 Continual Learning of LMs on Long Task Sequences
LoRA (Hu et al., 2021) $r\,(d_{\text{in}}+d_{\text{out}})$
@@
MoE-Adapter (2 experts/task) (Yu et al., 2024) $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
CoDyRA (Lu et al., 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
O-LoRA (Wang et al., 2023) $r\,(d_{\text{in}}+d_{\text{out}})$
LB-CL (Qiao & Mahdavi, 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
MoRA $rd_{\text{in}}+kd_{\text{out}}$
A.3 Multi‐domain task incremental learning.
We evaluate MoRA in the few‐shot MTIL setting (Table 11) under the same protocols as (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024). Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
@@
Table 11: Comparisons on 5-shot MTIL setting. Following the same protocol as in (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
   MoE-Adapter (Yu et al., 2024) 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
   RAIL-Primal (Xu et al., 2024) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   CoDyRA (Lu et al., 2024) 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
@@
   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
   MoE-Adapter (Yu et al., 2024) 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
   RAIL-Primal (Xu et al., 2024) 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
   CoDyRA (Lu et al., 2024) 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
@@
   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
   MoE-Adapter (Yu et al., 2024) 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
   RAIL-Primal (Xu et al., 2024) 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
   CoDyRA (Lu et al., 2024) 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
A.4 More Comparison on HumanEval for Code Generation
To further evaluate code generation performance, we compare MoRA against LoRI-D and LoRI-S (Zhang et al., 2025) on the HumanEval benchmark (Table 12). MoRA consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.

### v3

@@ 15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
Table 7: The six task‐sequence orders used in our continual learning experiments. Sequences 1–3 follow the standard CL benchmarks employed in prior work. Sequences 4–6 extend to longer 15‐task streams, as introduced in [65].
@@
A.2 Evaluation Metrics
To strictly evaluate the plasticity and stability of our method, we employ the following metrics. Let $N$ denote the total number of learned tasks, and $A_{i,j}$ denote the performance (e.g., accuracy, ROUGE, or similarity score) on task $j$ after training on task $i$.
CLIP Continual Learning Metrics.
For the CLIP experiments (MTIL and X-TAIL), we report the following metrics as in [105, 96, 91]:
  • Transfer Accuracy (Transfer): indicates the model’s zero-shot performance on future domains before they are learned, measures the extent to which the pre-trained zero-shot ability is preserved (or improved) throughout incremental learning.
  • Average Accuracy (Average): indicates the average accuracy of all learning steps across all domains. It captures the comprehensive performance stability throughout the entire incremental training process.
  • Last Accuracy (Last): represent the model’s performance on all seen domains after the training is fully completed.
LLM Continual Learning Metrics.
For the TRACE benchmark, we adopt the standard metrics as in [81]:
  • Overall Performance (OP): The average performance across all learned tasks at the end of training.
    $\text{OP}=\frac{1}{N}\sum_{j=1}^{N}A_{N,j},$ (9)
    where $A_{i,j}$ represents the performance on task $j$ after learning task $i$.
  • Backward Transfer (BWT): Measures the average performance degradation (forgetting) of previous tasks $j<N$ after learning new tasks. A lower BWT indicates less forgetting (better stability).
    $\text{BWT}=\frac{1}{N-1}\sum_{j=1}^{N-1}(A_{j,j}-A_{N,j}).$ (10)
A.3 Continual Learning of LMs on the Standard-CL benchmark.
Table 8 reports results across three task orderings: MoRAM consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound. Unlike O-LoRA [80] and LB-CL [61], which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRAM needs no extra regularization. By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRAM lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns. We further evaluate MoRAM on LLaMA2-7B [72] under the same continual-learning setup (Table 10 in the Appendix), MoRAM also outperforms O-LoRA by 2.3% averaged over 3 task orders.
A.4 Continual Learning of LMs on Long Task Sequences
In Table 8, we evaluate MoRAM on standard CL benchmarks [65]. In Table 9, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in [80, 61]. Consistent with the findings in Table 8, MoRAM outperforms previous methods in terms of averaged performance across three task orders, and largely closes the gap to multi-task learning. Two key design choices drive this robust performance in the long‐sequence regime. First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRAM encourages each rank to specialize on a narrow subspace of the data manifold. At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference. Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows. Coupling with our proposed rank pruning, these mechanisms ensure that MoRAM continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
Table 8: Summary of results on standard CL benchmarks with T5-large. We report the averaged accuracy after training on the last task across three task orderings.
Standard CL Benchmark
Method Order-1 Order-2 Order-3 Avg.
MTL 80.0
SeqFT 18.9 24.9 41.7 28.5
SeqLoRA 44.6 32.7 53.7 43.7
IncLoRA 66.0 64.9 68.3 66.4
Replay 55.2 56.9 61.3 57.8
EWC 48.7 47.7 54.5 50.3
LwF 54.4 53.1 49.6 52.3
L2P 60.3 61.7 61.1 60.7
LFPT5 67.6 72.6 77.9 72.7
InfLoRA 75.2 75.4 75.8 75.5
O-LoRA 75.4 75.7 76.3 75.8
LB-CL 76.9 76.5 76.8 76.7
MoRAM 77.4 77.5 77.9 77.6
@@ A.4 Continual Learning of LMs on Long Task Sequences
Table 9: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from [61].
@@
MoRAM 77.8 78.0 79.3 78.4
Table 11: Comparisons of trainable parameters for each pre-trained weight matrix during continual learning of each task. MoE-LoRA and MoE-Adapter trains additional router module with LoRA experts. In MoRAM, $k$ denotes the rank activation budget (with $k\leq r$).
@@
LoRA [33] $r\,(d_{\text{in}}+d_{\text{out}})$
@@
MoE-Adapter (2 experts/task) [96] $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
CoDyRA [50] $r\,(d_{\text{in}}+d_{\text{out}})+r$
O-LoRA [80] $r\,(d_{\text{in}}+d_{\text{out}})$
LB-CL [61] $r\,(d_{\text{in}}+d_{\text{out}})+r$
MoRAM $rd_{\text{in}}+kd_{\text{out}}$
A.5 Multi‐domain task incremental learning.
We evaluate MoRAM in the few‐shot MTIL setting (Table 12) under the same protocols as [96, 91, 50]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
@@
Table 12: Comparisons on 5-shot MTIL setting. Following the same protocol as in [96, 91, 50].
    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot [63] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [44] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [20] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [86] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [105] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
   MoE-Adapter [96] 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
   RAIL-Primal [91] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   CoDyRA [50] 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
   MoRAM 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
@@
   LwF [44] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [20] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [86] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [105] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
   MoE-Adapter [96] 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
   RAIL-Primal [91] 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
   CoDyRA [50] 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
   MoRAM 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
@@
   LwF [44] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [20] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [86] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [105] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
   MoE-Adapter [96] 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
   RAIL-Primal [91] 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
   CoDyRA [50] 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
   MoRAM 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
A.6 More Comparisons on CL on LMs
Results on SuperNI benchmark. To provide a complete comparison, we additionally evaluate MoRAM against SAPT [103] on the SuperNI benchmark [83] using the T5-Large backbone (Table 13). We report results in both replay-based and replay-free settings. Under the replay-based protocol introduced by SAPT, where pseudo-samples are generated using a trained generative model, MoRAM attains a stronger stability–plasticity trade-off (AP: 51.79% vs. 51.54%; FT: 0.73% vs. 0.91%). In the replay-free setting, which is the primary use case for MoRAM, our method achieves state-of-the-art performance (AP: 39.62%) and substantially outperforms parameter-efficient baselines such as O-LoRA (26.37%) and L2P (15.18%). These results highlight the complementary design choices. SAPT improves replay efficiency through shared attention prompts, while MoRAM relies on architectural isolation via sparse rank-1 experts. As a result, MoRAM remains effective without a memory buffer, yet can also incorporate replay to further improve performance beyond SAPT.
Discussions on performance robustness. To assess order robustness, we incorporate the OPD metric proposed by [95], which measures a model’s sensitivity to the sequence of arriving tasks. Following standard practice for evaluating global performance stability, we compute the standard deviation of the final average accuracy over the $K$ task orders considered. In Table 14, using results on the Standard CL Benchmark with three distinct task orders, MoRAM demonstrates substantially improved robustness to task ordering. The disparity across orders is 0.26 for MoRAM, which is approximately half of that of O-LoRA (0.46). This indicates that the Self-Activated Sparse Mixture mechanism effectively reduces task interference and mitigates the unidirectional knowledge transfer effects identified in prior work.
Table 13: Overall results on the SuperNI Benchmark using the T5-Large backbone. We report Average Performance (AP, $\uparrow$) and Forgetting (FT, $\downarrow$). The best results for the stability-plasticity trade-off are highlighted in bold for methods with and without memory replay, respectively.
Methods Replay SuperNI
AP FT
Replay-Based Methods
Replay 35.37 16.92
SAPT 51.54 0.91
MoRAM 51.79 0.73
Replay-Free Methods
L2P 15.18 3.65
IncLoRA 12.33 41.93
C-LoRA 22.69 24.25
O-LoRA 26.37 19.15
MoRAM 39.62 5.74
Table 14: Order Robustness Analysis on standard CL benchmark with T5-Large. We report the accuracy for each order and the average accuracy $\pm$ the standard deviation.
Method Order 1 Order 2 Order 3 Avg $\pm$ Std
SeqFT 18.9 24.9 41.7 28.5 $\pm$ 11.82
L2P 60.3 61.7 61.1 61.0 $\pm$ 0.70
LFPT5 67.6 72.6 77.9 72.7 $\pm$ 5.15
O-LoRA 75.4 75.7 76.3 75.8 $\pm$ 0.46
MoRAM 77.4 77.5 77.9 77.6 $\pm$ 0.26
A.7 More Comparison on HumanEval for Code Generation
To further evaluate code generation performance, we compare MoRAM against LoRI-D and LoRI-S [99] on the HumanEval benchmark (Table 15). MoRAM consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.